New Gadgets, Inc., currently pays no dividend but is expected to pay its first annual dividend of $4.85 per share exactly 6 years from today. After that, the dividends are expected to grow at 3.4 percent forever. If the required return is 11.2 percent, what is the price of the stock today?  
Answer #1

The stock that pays no dividends for 5 years. Once the stock begins paying dividends, it will have a constant growth rate of dividends.

Now, we will use we the dividend in Year 6 and find the stock price in Year 5.

Pt = [Dt × (1 + g)] / (R − g)

P5 = D6 / (R − g) = $4.85 / (0.112 − 0.034) = $62.179

The price of the stock today is simply the PV of the stock price in the future. We simply discount the future stock price at the required return. The price of the stock today will be:

P0 = $62.179 / 1.112^5

P0 = $36.57